OpenStudy (bekkah323):

Find an equation of variation in which y varies inversely as x and y = 100 when x=25

jimthompson5910 (jim_thompson5910):

y varies inversely as x so y = k/x

jimthompson5910 (jim_thompson5910):

it tells us that when x = 25, y = 100 so y = k/x y = k/25 .. plug in x = 25 100 = k/25 .. plug in y = 100 now solve for k

OpenStudy (bekkah323):

i got 2500 but i feel like thats wrong?

jimthompson5910 (jim_thompson5910):

2500 is correct, so the equation is y = 2500/x

jimthompson5910 (jim_thompson5910):

notice if you plugged in x = 25, you would get y = 100

OpenStudy (bekkah323):

ahhh ok thank you

jimthompson5910 (jim_thompson5910):

you're welcome